# Three-dimensional quantum criticality

### v2

@@ Abstract
Quantum critical phenomena are widely studied across various materials families, from high temperature superconductors to magnetic insulators. They occur when a thermodynamic phase transition is suppressed to zero temperature as a function of some tuning parameter such as pressure or magnetic field. This generally yields a point of instability – a so-called quantum critical point – at which the phase transition is driven exclusively by quantum mechanical fluctuations. Here we show that the heavy fermion metamagnet UTe2 possesses a quantum phase transition at extreme magnetic field strengths of over 70 T. Surprisingly, rather than terminating at one singular point, we find that the phase boundary is sensitive to magnetic field components in each of the three cartesian axes of magnetic field space. This results in the three-dimensional transition surface being bounded by a continuous ring of quantum critical points, the locus of which forms an extended line of quantum criticality – a novel form of quantum critical phase boundary. Within this quantum critical line sits an intensely field–resilient superconducting state in a striking toroidal shape, persisting to fields over 70 T. We model our data by a phenomenological free energy expansion, and show how a three-dimensional quantum critical phase boundary – rather than a more conventional singular point of instability – anchors the remarkable high magnetic field phase landscape of UTe2.
@@
At absolute zero temperature classical phases of matter are completely frozen in place. By contrast, in emergent electronic phases quantum zero-point motion allows for fluctuations to persist down to temperature $T=$ 0 K. For example, when the critical end point of a first-order transition is suppressed to zero temperature, a quantum phase transition [1] occurs at a quantum critical end point (QCEP) [2]. The enhancement of magnetic or density fluctuations close to such a transition can provide the pairing glue for unconventional superconducting phases to condense [3]. For example, $d$-wave superconductivity often accompanies quantum phase transitions involving antiferromagnetism [4, 5], whereas proximity to ferromagnetic ordering may promote the formation of exotic $p$-wave superconductive phases [6, 7, 8]. In a magnetic field the spin degree of freedom of $p$-wave superconductors can yield complex multi-component superconducting phase diagrams [9, 10]. In some rare cases, such as the ferromagnetic superconductor URhGe, the application of a magnetic field H has been found to induce superconductivity close to a high-$H$ QCEP [11, 12, 13].
@@
The paramagnetic metal UTe2 is another intriguing superconducting material. UTe2 possesses a remarkably rich phase diagram, with up to six distinct superconductive phases having been reported under various conditions of magnetic field tilt angle, temperature and pressure [14, 15, 16, 17, 18, 19]. At $\mu_{0}H$ = 0 T a superconducting ground state condenses at $T_{\text{c}}=$ 2.1 K [20, 21, 15, 22]. Despite the absence of ferromagnetism [23], this superconductive state exhibits multiple characteristics of odd-parity $p$-wave pairing [20, 15, 24] analogous to its ferromagnetic uranium-based cousins UGe2 [6], URhGe [7] and UCoGe [8]. The rich multidimensional phase space of UTe2 is further complicated by the presence of charge and pair density wave orderings that have been reported [25, 26] at low-$H$, while at high-$H$ a first-order metamagnetic (MM) transition to a field-polarised paramagnetic (PPM) state occurs for H $\parallel b$, the hard magnetic direction [27].
High field superconductivity in UTe2
@@
Of this menagerie of exotic emergent phases, the focus of the present study is the highest-$H$ superconducting phase found to date in UTe2, which is located at extremely high field strengths starting at $\mu_{0}H\approx$ 40 T and persisting to $\mu_{0}H\gtrapprox$ 70 T. Prior studies have reported this high-$H$ superconductivity to be entirely confined inside the PPM state, nestling against the PPM phase boundary within a narrow angular range for H tilted in the crystallographic $b-c$ plane [18, 28, 16, 29, 30, 31]. Recent measurements have also confirmed its existence for a component of H tilted away from the $b-c$ plane towards the $a$-axis [32] – but again, this exotic high-$H$ superconductivity was found exclusively inside the PPM state. The MM transition into the PPM state has been reported to be strongly first-order at low temperatures for the orientations measured so far [27, 29, 33, 16, 31, 30, 34].
[…]
In combination, our high-$H$ measurements (i) identify that an extended line of QCEPs bounds the surface of first-order MM transitions in three-dimensional H-space at the remarkably high magnetic field scale of over 70 T, (ii) anchor the field–resilient superconductivity in UTe2 close to this boundary, and (iii) reveal that in a narrow angular range close to the QCL the onset of field–resilient superconductivity extends down to fields that lie far below the MM transition field.
Quantum criticality at intense magnetic fields
The highest-field superconducting state in UTe2 is anchored to a quantum critical line associated with a field-induced metamagnetic transition (Figure 1). The key finding of our study is that rotating the orientation of H away from H $\parallel b$, for all directions, the critical end point of the MM transition is suppressed in temperature, eventually reaching 0 K. This thereby forms an extended line (the QCL, labelled in Fig. 1b) composed of QCEPs, demarcating a quantum phase transition at the termination of the PPM phase. Within the QCL sits the high-$H$ superconducting state (labelled SC3), forming a toroidal shape when plotted on Cartesian field axes (Fig. 1a).
We map this evolution of the UTe2 MM transition through contactless resistivity measurements by the proximity detector oscillator (PDO) technique (see Appendix A for details) in pulsed magnetic fields. In Fig. 2a we plot the PDO frequency shift for rotations in the plane that connects the $a$-axis ($\theta_{a}=$ 90°$\degree$, $\theta_{b-c}=$ 0°$\degree$) to the $b-c$ plane at $\theta_{b-c}=$ 17°$\degree$, $\theta_{a}=$ 0°$\degree$; panel (c) shows similar measurements for rotations between the high symmetry directions $b$ and $ac$, i.e. [010]–[101], and from $b$ to $a$, [010]–[100]. For H aligned close to the $b$-axis (i.e. at low values of $\theta_{a}$, $\theta_{b-a}$ and $\theta_{b-ac}$) at low temperatures the MM transition in the contactless resistivity is sharp and first order-like in character, consistent with prior studies [18, 36, 38, 39, 30, 40].
Tilting H away from $b$ increases the magnetic field strength $H_{m}$ at which the transition occurs. As $H_{m}$ increases, the sharpness of the transition smoothly decreases. This softening of the transition with rotation angle at low temperatures is very similar to the behavior for H $\parallel b$ at elevated temperatures plotted in Fig. 2d. At $T=$ 1.5 K the MM jump in PDO frequency is very sharp, exhibiting a large discontinuity in the derivative of the signal with respect to field. The critical end point of UTe2’s MM transition for H $\parallel b$ is reached at $T\approx$ 10 K; above this temperature the transition evolves into a broad crossover [16]. In our PDO measurements at $T=$ 9.0 K for H $\parallel b$ there is still a sharp jump at the transition – by contrast, close to the QCL at $\theta_{b-ac}=$ 42°$\degree$ the transition is markedly softer at $T=$ 1.5 K than it was at 9.0 K for H $\parallel b$.
This suppression of the size of the MM transition is even more apparent in our highest field measurements. In Fig. 2 we plot rotations from $b$ towards $a$ and from $\theta_{b-c}=$ 17°$\degree$ towards $a$, while in Fig. 3 we show the evolution for $\theta_{b-c}=$ 7°$\degree$ towards $a$ and from $b$ towards $c$. In Fig. 3c we show how these measurements enabled us to map the shape of the QCL in three-dimensional H-space. For each rotational plane probed by our experiments, we identify the location of that plane’s QCEP as the angle at which the size of the discontinuity in the derivative of the PDO signal tends to zero, thereby indicating the suppression of the first-order MM phase boundary towards zero temperature. These locations are marked with green stars in Fig. 1 and Fig. 3c. The locus of these QCEPs forms the QCL in three-dimensional H-space.
Condensation of SC3 before the metamagnetic transition
@@
In the ferromagnetic superconductor URhGe, H–induced superconductivity is found in close proximity to a QCEP at $\mu_{0}H\approx$ 10 T, identified by the suppression of a first-order MM transition to zero temperature [12]. Our finding of a similar phenomenon in UTe2 – but in this case exhibiting the remarkable geometric property of quantum criticality stretched along an extended line in H-space at the intense field strengths of 50 T $<\mu_{0}H\lessapprox$ 80 T – motivates a closer examination of SC3 close to the QCL, to which we now turn.
@@
Figure 4 shows contacted resistivity measurements by the standard four-terminal technique for a high quality UTe2 specimen (with a residual resistivity ratio of 420 – see Fig. 9). Up to the maximum applied field of the magnet system utilized for this suite of measurements, of 41.5 T, for incremental magnetic field tilt angles rotating from $b$ towards $ac$ at $T=$ 0.4 K we suddenly access the SC3 state at $\theta_{b-ac}=$ 17.5°$\degree$. In Fig. 4b we define the purple SC3 region by the observation of zero resistivity $\rho$ (in the low current limit). Interestingly, we observe a broadening of the onset of the SC3 phase for successively higher rotation angles $\theta_{b-ac}$, identified as the turning over of $\rho(H)$ from a positive to a negative gradient in $H$ and tracked by the dashed black line in Fig. 4a. This onset extends down to a minimum magnetic field of $\approx$ 30 T at $\theta_{b-ac}\approx$ 34°$\degree$, before progressing back up to higher fields for further inclination in the $b-ac$ plane.
@@
This wide magnetic field domain of negatively sloped magnetoresistance exhibits a highly non-linear relationship between excitation current $I$ and voltage $V$. Whereas standard Ohmic behavior is observed for all measured temperatures at 25 T when the sample is in the normal state, at 41 T a strong deviation is observed at low temperatures (Fig. 4e,f), indicating the likely presence of exotic vortex dynamics [41]. In Figure 5 we plot the exponent $\alpha$ fitted to $V\propto I^{\alpha}$ to illustrate this departure from Ohmic behavior at low temperatures for $\mu_{0}H\gtrapprox$ 30 T. This non-linearity between $I$ and $V$ clearly extends down far below $H_{m}$. We therefore identify a broad range of the UTe2 high field phase diagram, for H tilted close to the QCL, to be occupied by the onset of the SC3 phase.
@@
For H $\parallel b$, the first-order MM transition to the PPM is observed at $\mu_{0}H_{m}\approx$ 34.5 T in the zero temperature limit [27, 29, 42, 36]. As discussed above, on warming the critical end point is crossed at $T\approx$ 10 K and the transition then evolves into a broad crossover, characterized by e.g. a maximum in $\rho(H)$, as the PPM gradually melts and $H_{m}$ decreases until $T\approx$ 30 K where $\mu_{0}H_{m}$ reaches 0 T [15, 16, 27, 29, 34]. Accordingly, we identify the location of $H_{m}$ for each temperature as the field strength at which $\nicefrac{{\partial\rho}}{{\partial H}}=0$, as shown in Figure 6. We schematically plot the corresponding phase landscapes for H $\parallel b$ (i.e. $\theta_{b-ac}=0\degree$) and $\theta_{b-ac}=35\degree$ in Fig. 4c. Strikingly, upon warming above $T_{c}^{SC3}$ at $\theta_{b-ac}=35\degree$, we do not observe a signature of the PPM state in our accessible field range of $\mu_{0}H\leq$ 41.5 T until $T\approx 15$ K (Fig. 6).
@@
Furthermore, we performed a series of rotations in the $b-ac$ plane at 6 K – to suppress all superconducting states and therefore solely probe the PPM phase – to map out the boundary of the MM transition in this plane, which we plot as red circles in Fig. 4b (see Fig. 10 for raw data). The MM transition clearly bisects a substantive region of SC3. Therefore, our measurements have uncovered that a portion of the SC3 phase – and its broad onset region – extends considerably outside the PPM state in the $b-ac$ plane as H is tilted close to the QCL, as we depict in Figure 7. This is in stark contrast to prior measurements of SC3 in the $b-c$ plane, comparatively far from the QCL, that observed SC3 exclusively within the PPM host-state [18, 28, 16, 29, 30, 31, 40].
Landau theory of quantum criticality in three dimensions

### v3

# A quantum critical line bounds the high field metamagnetic transition surface in UTe2

@@ Abstract
Quantum critical phenomena are widely studied across various materials families, from high temperature superconductors to magnetic insulators. They occur when a thermodynamic phase transition is suppressed to zero temperature as a function of some tuning parameter such as pressure or magnetic field. This generally yields a point of instability – a so-called quantum critical point – at which the phase transition is driven exclusively by quantum fluctuations. Here we show that the heavy fermion metamagnet UTe2 possesses a quantum phase transition at extreme magnetic field strengths of over 70 T. Rather than terminating at one singular point, we find that the phase boundary is sensitive to magnetic field components in each of the three Cartesian axes of magnetic field space. This results in the three-dimensional transition surface being bounded by a continuous ring of quantum critical points, the locus of which forms an extended line of quantum criticality – a novel form of quantum critical phase boundary. Within this quantum critical line sits a magnetic field-induced superconducting state in a toroidal shape, which persists to fields over 70 T. We model our data by a phenomenological free energy expansion, and show how a three-dimensional quantum critical phase boundary – rather than a more conventional singular point of instability – anchors the remarkable high magnetic field phase landscape of UTe2.
I Introduction
@@
At absolute zero temperature classical phases of matter are completely frozen in place. By contrast, in emergent electronic phases fluctuations from quantum zero-point motion may persist down to temperature $T=$ 0 K. For example, when the critical end point (CEP) of a first-order transition is suppressed to zero temperature, a quantum phase transition [1] occurs at a quantum critical end point (QCEP) [2]. The enhancement of magnetic or density fluctuations close to a quantum critical point can provide the pairing glue for unconventional superconducting phases to condense [3]. For example, $d$-wave superconductivity often accompanies quantum phase transitions involving antiferromagnetism [4, 5], whereas proximity to ferromagnetic ordering may promote the formation of exotic $p$-wave superconductive phases [6, 7, 8]. In a magnetic field the spin degree of freedom of $p$-wave superconductors can yield complex multi-component superconducting phase diagrams [9, 10]. In some rare cases, such as the ferromagnetic superconductor URhGe, the application of a magnetic field H has been found to induce superconductivity close to a high-$H$ QCEP [11, 12, 13].
@@
The paramagnetic metal UTe2 possesses a remarkably rich phase diagram, with up to six distinct superconductive phases having been reported under various conditions of magnetic field tilt angle, temperature and pressure [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. At $\mu_{0}H$ = 0 T a superconducting ground state forms at $T_{\text{c}}=$ 2.1 K [24, 25, 26]. Despite the absence of ferromagnetism [27], this superconductive state exhibits multiple characteristics of odd-parity $p$-wave pairing [14, 17, 28] analogous to its ferromagnetic uranium-based cousins UGe2 [6], URhGe [7] and UCoGe [8]. The rich multidimensional phase space of UTe2 may further contain charge and pair density wave orderings that have been reported by surface-sensitive probes [29, 30] at low-$H$, while at high-$H$ a first-order metamagnetic (MM) transition to a field-polarized paramagnetic (PPM) state occurs for H $\parallel b$, the hard magnetic direction [31].
The focus of the present study is the highest-$H$ superconducting phase found to date in UTe2, which is located at extremely high field strengths starting from $\mu_{0}H\approx$ 40 T and extending up to $\mu_{0}H\gtrapprox$ 70 T. The identification of this superconducting state was first made by contacted and contactless conductivity measurements in pulsed magnetic fields [20]. It was subsequently corroborated by bulk thermodynamic [32] and Hall effect [33] measurements.
@@
Prior studies have reported this high-$H$ superconductivity to be entirely confined inside the PPM state, nestling against the PPM phase boundary within a narrow angular range for H tilted in the crystallographic $b-c$ plane [20, 34, 18, 35, 32, 33]. Recent measurements have also confirmed its existence for a component of H tilted away from the $b-c$ plane towards the $a$-axis [36] – but again, this exotic high-$H$ superconductivity was found exclusively inside the PPM state. The MM transition into the PPM state has been reported to be strongly first-order at low temperatures for the orientations measured so far [31, 35, 37, 38, 39, 18, 33, 32, 40, 41]. The shape and extent of the MM transition surface, which separates paramagnetic from field-polarized parts of the high field phase diagram, thereby plays an important role in understanding the drivers for the unusual high field superconductivity in UTe2. Here, we map out the transition surface, show that it is bounded by a critical line, and that the high-$H$ superconducting phase in turn is anchored against that line. These findings suggest a central role for metamagnetic fluctuations in the pairing interaction in UTe2 at high fields.
II Results
II.1 Quantum criticality in intense magnetic fields
We probed the high-$H$ phase diagram of UTe2 with contacted and contactless electrical conductivity measurements on high quality samples grown by a salt flux method [24, 42, 25, 26]. Figure 1 presents contactless magnetoconductance data obtained by the proximity detector oscillator (PDO) technique (see Appendix A for details) in pulsed magnetic fields. For H $\parallel b$ the MM transition of UTe2 occurs at a characteristic field strength of $\mu_{0}H_{m}\approx 34$ T. As the orientation of H is rotated away from $b$, $H_{m}$ increases [20, 33, 36].
This can be seen in Fig. 1c, in which we plot the PDO frequency shift $\Delta f_{\text{PDO}}$ for incremental rotations in the plane that connects the $a$-axis ($\theta_{a}=$ 90°$\degree$, $\theta_{b-c}=$ 0°$\degree$) to the $b-c$ plane at $\theta_{b-c}=$ 17°$\degree$, $\theta_{a}=$ 0°$\degree$, at temperature $T=$ 1.7 K. (See Fig. 11 for a description of this rotation sequence.) Close to $b$, the MM transition manifests as an abrupt increase in $\Delta f_{\text{PDO}}$, due to the dramatic increase in resistivity that occurs upon entering the PPM state [38]. As the orientation of H is tilted further away from the $b-c$ plane towards the $a$ axis, $H_{m}$ migrates to higher $H$. Notably, the abruptness of the transition gradually diminishes as $H_{m}$ increases: whereas at $\theta_{a}=$ 6°$\degree$ the jump in $\Delta f_{\text{PDO}}$ is very sharp, at $\theta_{a}=$ 17°$\degree$ the MM transition is markedly softer, and by $\theta_{a}=$ 20°$\degree$ the magnetoconductance curve is very smooth throughout, with no feature indicating the presence of metamagnetism occurring at this angle and temperature for $\mu_{0}H\leq$ 78 T.
This softening of the MM transition, as a function of rotation angle at fixed (low) temperature, is very similar to the behavior that occurs as a function of temperature at the fixed orientation of H $\parallel b$. In Fig. 1e we plot this evolution of $\Delta f_{\text{PDO}}$ as the CEP of the transition is crossed upon raising $T$. The CEP of UTe2’s MM transition for H $\parallel b$ is reached at $T\approx$ 10 K; above this temperature the transition evolves into a broad crossover [18]. For this orientation, at $T=$ 1.5 K the jump in $\Delta f_{\text{PDO}}$ is very sharp, exhibiting a large discontinuity in the derivative of the signal with respect to field (panel f). At $T=$ 9.0 K there is still a sharp discontinuity in $\nicefrac{{\partial f}}{{\partial H}}$; however, at 12 K $\nicefrac{{\partial f}}{{\partial H}}$ is much smoother, having crossed the CEP. By contrast, for H tilted 42°$\degree$ from [010] towards the [101] direction (labeled as $\theta_{b-ac}$, lower parts of panels e & f), for all temperatures down to 1.5 K the jump in $\nicefrac{{\partial f}}{{\partial H}}$ is notably smoother than that observed at 9.0 K for H $\parallel b$. This implies that the CEP is located at much lower $T$ for H tilted to $\theta_{b-ac}=42\degree$ compared to when H $\parallel b$. In combination, these measurements of the suppression of the MM transition – for variable tilt angles at fixed temperature, and at incremental temperatures for fixed orientation – demonstrate that the CEP of the UTe2 normal paramagnetic to polarized paramagnetic phase transition is suppressed towards a QCEP when H is rotated away from the $b$-axis. We sketch a schematic overview of this phenomenon in Fig. 1a,b.
In Figure 2 we track the evolution of UTe2’s MM transition through four additional rotation planes: from $b$ to $a$, $b$ to $ac$, $b$ to $c$ and at a fixed tilt of 7°$\degree$ from $b$ to $c$ subsequently rotated towards $a$. For each of these rotation planes we observe the same behavior as in Fig. 1c – namely, that as H is tilted away from $b$, the sharpness of the MM transition decreases as $H_{m}$ increases, characteristic of the CEP being suppressed towards a QCEP. This result is notable, when considering the geometrical properties of the PPM state’s phase boundary. Figs. 1 & 2 have shown the sensitivity of the MM transition to rotations of the orientation of H. These experiments were performed by incrementally rotating samples about an axis of rotation; hence we presented the data in terms of the angle of the relative orientation of H with respect to the crystallographic axes. Consider instead the vector H decomposed into its three orthogonal (Cartesian) bases – i.e. ($H_{a},H_{b},H_{c}$). In this three-dimensional magnetic field space, the MM transition surface that separates the normal paramagnetic and polarized paramagnetic phases of UTe2 is defined as the connected surface of points that is swept out by rotating H through each location of $\mathbf{H_{\textit{m}}}$  [^1] Note that $\mathbf{H_{\textit{m}}}$ is itself a function of ($H_{a},H_{b},H_{c}$).. In other words, the MM transition surface is the connected surface of points on which each point lies a distance |$\absolutevalue{\mathbf{H_{\textit{m}}}}$ | from the origin.
In Fig. 3a we schematically plot the MM transition surface (in orange) on Cartesian magnetic field axes, alongside each of the three (ambient pressure) superconductive phases of UTe2. We label these in ascending order of their upper critical fields as SC1, SC2 and SC3, respectively [18, 25]. For each rotational plane probed by our experiments, we identify the location of that plane’s QCEP as the orientation at which the size of the discontinuity in the derivative of the PDO signal tends to zero (marked with green stars), thereby indicating the suppression of the first-order MM phase boundary to zero temperature. The surface marking the first-order transition into the PPM state curves strongly along the $H_{a}$ direction and is nearly flat along the $H_{b}$ direction. Our findings show that it is bounded by an undulating quantum critical line, which is indicated in green in Fig. 3 – that we term a quantum critical line (QCL).
In Figure 4 we show how we constructed this mapping of the three-dimensional QCL of UTe2, by tracking the diminution of the MM transition in each measured rotation plane to obtain the location of that plane’s QCEP. We connect these experimentally determined QCEP locations with a smooth green line, to indicate the QCL that separates the polarized and normal paramagnetic states. Inside the region circumscribed by the QCL resides the highest-$H$ superconducting phase of UTe2 (SC3), which wraps around the $b$-axis in a toroidal shape (see Figs. 16 & 21 and ref. [36]).
II.2 Condensation of SC3 before the metamagnetic transition
@@
In the ferromagnetic superconductor URhGe, $H$-induced superconductivity is found in close proximity to a QCEP at $\mu_{0}H\approx$ 10 T, identified by the suppression of a first-order MM transition to zero temperature [12]. Our finding of a similar phenomenon in UTe2 – but in this case stretched along an extended line in three-dimensional H-space at the intense field strengths of 50 T $<\mu_{0}H\lessapprox$ 80 T – motivates a closer examination of SC3 close to the QCL, to which we now turn.
@@
Figure 5 shows contacted resistivity measurements by the standard four-terminal technique for a high quality UTe2 specimen (with a residual resistivity ratio of 420, see Fig. 10). Up to the maximum applied field of the magnet system utilized for this suite of measurements, of 41.5 T, for incremental magnetic field tilt angles rotating from $b$ towards $ac$ at $T=$ 0.4 K we suddenly access the SC3 state at $\theta_{b-ac}=$ 17.5°$\degree$. In Fig. 5b we define the purple SC3 region by the observation of zero resistivity $\rho$ (in the low current limit). Interestingly, we observe a broadening of the onset of the SC3 phase for successively higher rotation angles $\theta_{b-ac}$, identified as the turning over of $\rho(H)$ from a positive to a negative gradient in $H$. This is tracked by a dashed black line in Fig. 5a. The location of this maximum in $\rho(H)$, which appears to signify the onset of the SC3 state, extends down to a minimum magnetic field of $\approx$ 30 T at $\theta_{b-ac}\approx$ 34°$\degree$, before progressing back up to higher fields for further inclination in the $b-ac$ plane.
@@
This wide magnetic field domain of negatively sloped magnetoresistance exhibits a highly non-linear relationship between excitation current density $J$ and voltage $V$. Whereas standard Ohmic behavior is observed for all measured temperatures at 25 T when the sample is in the normal state, at 41 T a strong deviation is observed at low temperatures (Fig. 5e,f). This appears to indicate the occurrence of vortex motion [44, 45, 46], perhaps due to the presence of low pinning forces in this high quality sample. We note that exotic vortex properties have recently been reported for the SC1 phase at low $H$ [47]; whether the SC3 state also possesses anomalous vortex behavior is an open question. In Figure 6 we plot the exponent $\alpha$ fitted to $V\propto J^{\alpha}$ to illustrate the departure from Ohmic behavior at low temperatures for $\mu_{0}H\gtrapprox$ 30 T. This non-linearity between $J$ and $V$ clearly extends down far below $H_{m}$. We therefore identify a broad range of the UTe2 high field phase diagram, for H tilted close to the QCL, to be occupied by the onset of the SC3 phase.
@@
For H $\parallel b$, the first-order MM transition to the PPM is observed at $\mu_{0}H_{m}\approx$ 34.5 T in the zero temperature limit [31, 35, 48, 25]. As discussed in Section II.1, on warming the CEP is crossed at $T\approx$ 10 K and the transition then evolves into a broad crossover, characterized by e.g. a maximum in $\rho(H)$, as the PPM gradually melts and $H_{m}$ decreases until $T\approx$ 30 K where $\mu_{0}H_{m}$ reaches 0 T [17, 18, 31, 35, 40]. Accordingly, we identify the location of $H_{m}$ for each temperature as the field strength at which $\nicefrac{{\partial\rho}}{{\partial H}}=0$, as shown in Figure 7. We schematically plot the corresponding phase landscapes for H $\parallel b$ (i.e. $\theta_{b-ac}=0\degree$) and $\theta_{b-ac}=35\degree$ in Fig. 5c. Strikingly, upon warming above $T_{c}^{SC3}$ at $\theta_{b-ac}=35\degree$, we do not observe a signature of the PPM state in our accessible field range of $\mu_{0}H\leq$ 41.5 T until $T\approx 15$ K (Fig. 7).
@@
Furthermore, we performed a series of rotations in the $b-ac$ plane at 6 K – to suppress all superconducting states and therefore solely probe the PPM phase – to map out the boundary of the MM transition in this plane, which we plot as red circles in Fig. 5b (see Fig. 12 for raw data). The MM transition clearly bisects a substantive region of the SC3 domain (as determined by the observation of zero $\rho$ in the limit of low $J$ and $T$). Therefore, our measurements have uncovered that a portion of the SC3 phase – and its broad onset region – extends considerably outside the PPM state in the $b-ac$ plane as H is tilted close to the QCL (Fig. 8). This is in stark contrast to prior measurements of SC3 in the $b-c$ plane, comparatively far from the QCL, that observed SC3 exclusively within the PPM host-state [20, 34, 18, 35, 32, 33, 49].
II.3 Landau theory of metamagnetic transitions in three dimensions